# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v3

@@ Abstract.
Quasi-cyclic codes have been recently employed in the constructions of quantum error-correcting codes. In this paper, we propose a construction of infinite families of quasi-cyclic codes which are self-orthogonal with respect to the Euclidean and Hermitian inner products. In particular, their dimension and a lower bound for their minimum distance are computed using their constituent codes defined over field extensions of $\mathbb{F}_{q}$. We also show that the lower bound for the minimum distance satisfies the square-root-like lower bound and also show how self-dual quasi-cyclic codes can arise from our construction. Using the CSS construction, we show the existence of quantum error-correcting codes with good parameters.
Key words and phrases:
quasi-cyclic codes, self-dual codes, self-orthogonal codes, CSS construction, quantum error correcting codes
2020 Mathematics Subject Classification:
94B60
1. Introduction
Quantum error correction plays a central role in the development of quantum computing architectures. With the rapid advancement of technologies involving quantum computation and communication, there is a growing demand for effective quantum error-correcting codes (QECCs). These codes are essential for protecting quantum information from noise, which can arise from various types of quantum channels. In the celebrated work of Calderbank, Shor, and Steane [Calderbank-Shor1996, Steane1996], the authors proposed the CSS construction that utilizes the duality of certain classical error-correcting codes to construct quantum error-correcting codes. This method is commonly used to construct QECCs over finite fields [Ketkar-Klappenecker-Kumar-Sarvepalli2006]. QECCs from the CSS construction are most often constructed using self-orthogonal error-correcting codes and its dual, where self-orthogonal codes have applications in lattice theory [W98], row-self-orthogonal matrices [Massey92, Massey98], linear complementary dual codes [Massey98], and many others. Beyond self-orthogonality, the notions of dual-containing and self-dual codes are also significant. Self-dual codes notably exhibit deep connections with group theory [Hoehn00, JLLX13], lattices [SloaneChapter99], and design theory [Sole90], and are recognized for being asymptotically good [MACWILLIAMS1972153]. Because of these applications, it is a worthwhile endeavor to determine the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes.
The focus of this paper is on constructing quasi-cyclic (QC) error-correcting codes that are self-orthogonal and self-dual under the Euclidean and Hermitian inner products. Quasi-cyclic codes are a natural generalization of cyclic codes and are well-studied in classical coding theory literature. Quasi-cyclic codes have a rich algebraic structure [seguin, cem, lingsole] and have strong connections with convolutional codes [SolomonTilborg]. Moreover, there exist QC codes that are asymptotically good [longcodes]. In 2022, the authors in [RB] found an upper bound for the locality of QC locally recoverable codes using generator matrices in systematic form where the first block is the identity matrix and the remaining blocks are comprised of smaller circulant matrices. More recently in  [BastosAlvarezFloresSalerno], a more concise way of finding an upper bound on the locality of these codes was studied, and a construction of optimal locally recoverable QC codes was presented.
@@ 1. Introduction
Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in [Abdukhalikov-Bag-Panario2023, GalindoHernando, benjwal], the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. Consequently, as these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
@@
In this paper, we propose a construction of infinite families of self-orthogonal and self-dual QC codes using their constituent code representation described in [cem]. We then describe how the codes in our construction yield the existence of QECCs with good parameters. The paper is organized as follows: Section 2 provides background results related to QC codes and their constituent codes. This section also includes the foundations of constructing QECCs from classical ones. In Section 3, we present the main results of this paper, namely, the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (or self-dual) with respect to the Euclidean inner product. We show that the self-orthogonal QC codes in this construction satisfy the square-root-like lower bound to their minimum distance. We also introduce the notion of Galois-closed codes to obtain QC codes that are self-orthogonal with respect to the Hermitian inner product. Lastly in Section 4, we show how some codes from our construction are used to obtain quantum error-correcting codes with the best known parameters so far. All computations throughout this paper have been done using SageMath or Mathematica.
2. Preliminaries
2.1. Quasi-Cyclic Codes and Their Representations
@@
(1) $d\leq n-k+1$
and the code is called a maximum distance separable code, or an MDS code, if equality in (1) is achieved. A quintessential example of a maximum distance separable code is the Reed-Solomon code described in [RS1958]. More generally, if we choose nonzero elements $v_{1},...,v_{n}\in\mathbb{F}_{q}$ and distinct elements $\alpha_{1},...,\alpha_{n}\in\mathbb{F}_{q}$, we define the Generalized Reed-Solomon Codes as
(2) $GRS_{n,k}(\mathbf{\alpha},\mathbf{v})=\{(v_{1}f(\alpha_{1}),...,v_{n}f(\alpha_{n}))\mid f(x)\in\mathbb{F}_{q}[x]_{k}\}$
@@
(3) $\textbf{c}=\begin{pmatrix}c_{00}&c_{01}&...&c_{0,\ell-1}\\ c_{10}&c_{11}&...&c_{1,\ell-1}\\ \vdots&\vdots&\ddots&\vdots\\ c_{m-1,0}&c_{m-1,1}&...&c_{m-1,\ell-1}\end{pmatrix}$
where the action of the operator $T^{\ell}$ is equivalent to a row shift in (3). We can also “flatten the array” by its rows to obtain the vector
(4) $\mathbf{c}=(c_{0,0},c_{0,1},\ldots,c_{0,\ell-1},c_{1,0},c_{1,1},\ldots,c_{1,\ell-1},\ldots,c_{m-1,0},c_{m-1,1},\ldots,c_{m-1,\ell-1})\in\mathbb{F}_{q}^{m\ell}.$
Moreover, as in [lingsole], we consider the ring $R_{q,m}^{\ell}:=\left(\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle\right)^{\ell}$ where $\operatorname{gcd}(m,q)=1$. To any codeword c as in (3), we associate an element of $R_{q,m}^{\ell}$:
(5) $\vec{\textbf{c}}(x):=\left(c_{0}(x),c_{1}(x),...,c_{\ell-1}(x)\right)\in R_{q,m}^{\ell}$
@@
(6) $c_{j}(x):=c_{0j}+c_{1j}x+c_{2j}x^{2}+\cdots+c_{m-1,j}x^{m-1}\in R_{q,m}.$
By Lemma 3.1 in [lingsole], the map $\phi:\mathbb{F}_{q}^{m\ell}\to R_{q,m}^{\ell}$ defined by $\phi(\textbf{c})=\vec{\textbf{c}}(x)$ is an $R$-module isomorphism. Hence, any QC code over $\mathbb{F}_{q}$ of length $m\ell$ and index $\ell$ can be viewed as an $R$-submodule of $R_{q,m}^{\ell}$. Throughout this paper, we will move freely between the representations in (4) and (5).
Given a polynomial $f(x)=f_{n}x^{n}+f_{n-1}x^{n-1}+\cdots+f_{0}\in\mathbb{F}_{q}[x]$, its reciprocal polynomial, $f^{*}$, is given by $f^{*}(x)=x^{n}f(x^{-1})$. As seen in [lingsole], we may decompose $x^{m}-1\in\mathbb{F}_{q}[x]$ into irreducible factors
(7) $x^{m}-1=\delta\prod_{i=1}^{t}g_{i}(x)g_{i}^{*}(x)\prod_{i=1}^{s}f_{i}(x),$
where $\delta$ is a non-zero element in $\mathbb{F}_{q}$, $f_{i}(x)$ is a self-reciprocal polynomial for each $1\leq i\leq s$, and $g_{i}^{*}(x)$ is the reciprocal polynomial of $g_{i}(x)$, where $g_{i}(x)\neq g_{i}^{*}(x)$, for each $1\leq i\leq t$. With this decomposition and via the Chinese remainder theorem (CRT), the ring $R_{q,m}$ may be written as
(8) $R_{q,m}=\frac{\mathbb{F}_{q}[x]}{\langle x^{m}-1\rangle}\cong\Bigg(\bigoplus_{j=1}^{t}\bigg(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\bigg)\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg),$
where we have suppressed the explicit dependence on $x$ in the $g_{j}$ and $f_{i}$. Let $\alpha$ be an $m$-th primitive root of the unity in an extension of $\mathbb{F}_{q}$ and let
$F_{i}:=\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\cong\mathbb{F}_{q}(\alpha^{u_{i}}),\quad G^{\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g_{i}\rangle}\cong\mathbb{F}_{q}(\alpha^{v_{j}}),\quad\text{and}\quad G^{\prime\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\cong\mathbb{F}_{q}(\alpha^{v^{*}_{j}}).$
@@
(9) $R_{q,m}^{\ell}\cong\Bigg(\bigoplus_{j=1}^{t}(G^{\prime}_{j}\oplus G^{\prime\prime}_{j})^{\ell}\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}F_{i}^{\ell}\bigg).$
@@
Define codes $C_{i}\subset F_{i}^{\ell}$, $C^{\prime}_{j}\subset G_{j}^{{}^{\prime}\ell}$, and $C^{\prime\prime}_{j}\subset G_{j}^{{}^{\prime\prime}\ell}$ for $1\leq i\leq s$ and $1\leq j\leq t$. We call these codes the constituent codes of the $R_{q,m}$-QC code $\mathcal{C}\subset R_{q,m}^{\ell}$ which has the following CRT decomposition:
(10) $C\cong\Bigg(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus C_{j}^{\prime\prime})\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}C_{i}\bigg).$
@@
The following lemma gives an easy way to compute the dimension of a QC code $\mathcal{C}$ from its constituent codes. We slightly modified the notation of the following result from [BastosAlvarezFloresSalerno] to align with the notation in this paper.
Lemma 2.1.
@@
[BastosAlvarezFloresSalerno, Lemma 2] Let $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ be an $m\ell,k]$ QC code and let $C_{j}\subseteq\mathbb{F}_{q}^{\ell}(\alpha^{u_{j}})\cong\left(\frac{\mathbb{F}_{q}[x]}{\langle b_{j}\rangle}\right)^{\ell}$ be its respective constituent $\ell,k_{j}]$ codes. Suppose that the nontrivial constituent codes of $\mathcal{C}$ are given by $C_{1},\ldots,C_{h}$. Then,
(11) $k=\sum_{j=1}^{h}k_{j}\deg(b_{j}).$

### v4

@@ Abstract
Quasi-cyclic codes have been recently employed in the constructions of quantum error-correcting codes. In this paper, we propose a construction of infinite families of quasi-cyclic codes over $\mathbb{F}_{q}$ which are self-orthogonal with respect to the Euclidean and Hermitian inner products. In particular, their dimension and a lower bound for their minimum distance are computed using their constituent codes defined over field extensions of $\mathbb{F}_{q}$. We also show that the lower bound for the minimum distance satisfies the square-root-like lower bound and also show how dual-containing and self-dual quasi-cyclic codes can arise from our construction. Using the CSS construction, we show the existence of quantum error-correcting codes with good parameters.
keywords:
[…]
1 Introduction
In the study of infinite families of codes, the growth rate of the minimum distance with respect to the code length is an important measure of performance. The square-root bound is a useful parameter for measuring how good a classical error-correcting code is; namely, given $n$ and $d$ the code length and the minimum distance of a code $C$, we say $C$ is a good code if $d\geq\sqrt{n}.$ This definition was extended to the square-root-like bound when there is a positive constant $c$ such that $d\geq c\sqrt{n}$. Codes satisfying square-root-like bounds are desirable as they exhibit nontrivial distance growth while maintaining strong algebraic structure. Such bounds guarantee that the error-correcting capability of the codes increases with the code length. Moreover, the square-root bound frequently arises as a natural benchmark for structured algebraic code families, where achieving linear distance growth is often difficult. The construction of infinite families of codes satisfying such bounds has attracted significant attention, and for instance, the various codes constructed in Chen2023FourIF; squarerootbound; 28fcd52cbe024304880942df5f3617ce; WLZZ25 satisfy the square-root-like bound.
Beyond minimum distance considerations, structural properties of codes such as self-orthogonality, dual-containment, and self-duality play important roles in coding theory. Self-orthogonal codes arise in several areas, including lattice theory W98, row-self-orthogonal matrices Massey92; Massey98, and linear complementary dual codes Massey98. Self-dual codes also exhibit deep connections with group theory Hoehn00; JLLX13, lattices SloaneChapter99, and design theory Sole90, and are known to be asymptotically good MACWILLIAMS1972153. The construction of codes that simultaneously satisfy the square-root-like bound together with certain duality conditions are also of interest. For instance, the authors in infiniteselfdualfamilies provide infinite families of (Euclidean and Hermitian) self-dual cyclic codes satisfying the square-root bound, and in goodsocycliccodes, the authors explore constructions of self-orthogonal cyclic codes whose minimum distances are larger than the square-root bound. Moreover, classical codes with self-orthogonal or dual-containing properties are useful in the construction of quantum error-correcting codes through celebrated work in Calderbank-Shor1996; Steane1996, known as the CSS construction. Because of these connections and applications, determining the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes is an important problem in coding theory.
According to infiniteselfdualfamilies, there are very few constructions of infinite families of self-dual codes whose minimum distances meet the square-root lower bound, which are depicted in that paper. To the best of our knowledge, the authors of infiniteselfdualfamilies in 2025 were the first to present infinite families of non-binary self-dual cyclic codes achieving a square-root-type lower bound on their minimum distances. Additionally, Calderbank provides a distinct construction of an infinite family of self-dual quasi-cyclic codes.
@@ 1 Introduction
Quasi-cyclic (QC) are have been suitable for constructing structured families of codes with desirable distance and duality properties. QC codes are a natural generalization of cyclic codes and are well-studied in the classical coding theory literature. They possess a rich algebraic structure seguin; cem; lingsole and have strong connections with convolutional codes SolomonTilborg. Moreover, there exist QC codes that are asymptotically good longcodes. Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in Abdukhalikov-Bag-Panario2023; GalindoHernando; benjwal, the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. As these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
@@
In this paper, we propose a construction of infinite families of self-orthogonal, dual-containing, and self-dual QC codes using their constituent code representation described in lingsole, and whose lower bound for their minimum distances are determined using their concatenated description described in jensen; OO24. We then show that the codes in our construction yield minimum distances that satisfy a square-root-like lower bound. We also explore the existence of QECCs with good parameters arising from these constructions. The paper is organized as follows: Section 2 provides background results related to QC codes and their constituent codes, and also includes the foundations of constructing QECCs from classical codes. In Section 3, we present the main results of this paper, namely the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (dual-containing, or self-dual) with respect to the Euclidean inner product. In Section 4, we show that minimum distances of such codes satisfy the square-root-like lower bound. We then introduce the notion of Galois closed codes in Section 5 to obtain QC codes that are self-orthogonal (dual-containing, or self-dual) with respect to the Hermitian inner product. Lastly, in Section 6, we show how some codes from our construction yield quantum error-correcting codes with good parameters using the CSS construction. All computations throughout this paper have been done using Magma, SageMath, and Mathematica.
uantum error correction plays a central role in the development of quantum computing architectures. With the rapid advancement of technologies involving quantum computation and communication, there is a growing demand for effective quantum error-correcting codes (QECCs). These codes are essential for protecting quantum information from noise, which can arise from various types of quantum channels. In the celebrated work of Calderbank, Shor, and Steane Calderbank-Shor1996; Steane1996, the authors proposed the CSS construction that utilizes the duality of certain classical error-correcting codes to construct quantum error-correcting codes. This method is commonly used to construct QECCs over finite fields Ketkar-Klappenecker-Kumar-Sarvepalli2006. QECCs from the CSS construction are most often constructed using self-orthogonal error-correcting codes and its dual, where self-orthogonal codes have applications in lattice theory W98, row-self-orthogonal matrices Massey92; Massey98, linear complementary dual codes Massey98, and many others. Beyond self-orthogonality, the notions of dual-containing and self-dual codes are also significant. Self-dual codes notably exhibit deep connections with group theory Hoehn00; JLLX13, lattices SloaneChapter99, and design theory Sole90, and are recognized for being asymptotically good MACWILLIAMS1972153. Because of these applications, it is a worthwhile endeavor to determine the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes.
The square-root bound is a useful parameter for measuring how good a classical error-correcting code is. Given $n$ and $d$ the code length and the minimum distance of a code $C$, we say $C$ is a good code if $d\geq\sqrt{n}.$ This definition was extended for square-root-like bound, when there is a positive constant $c$ so that $d\geq c\sqrt{n}$. Currently, the academic community have been focused on construct infinite families of codes satisfying the square-root bound, and their dualities. As exponents in this research area, we quote infiniteselfdualfamilies , where the authors provide infinity families of (Euclidean and Hermitian) self-dual cyclic codes satisfying the square-root. In goodsocycliccodes, the authors explore constructions of self-orthogonal cyclic codes whose their minimum distances are larger than the square-root bound.
The focus of this paper is on constructing quasi-cyclic (QC) error-correcting codes that are self-orthogonal, dual-containing and self-dual under the Euclidean and Hermitian inner products. Quasi-cyclic codes are a natural generalization of cyclic codes and are well-studied in classical coding theory literature. Quasi-cyclic codes have a rich algebraic structure seguin; cem; lingsole and have strong connections with convolutional codes SolomonTilborg. Moreover, there exist QC codes that are asymptotically good longcodes.
Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in Abdukhalikov-Bag-Panario2023; GalindoHernando; benjwal, the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. Consequently, as these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
In this paper, we propose a construction of infinite families of self-orthogonal and self-dual QC codes using their constituent code representation described in lingsole and their concatenated description described in jensen; OO24. We then describe how the codes in our construction yield codes whose minimum distances are also better than the square-root bound. We also explore the existence of QECCs with good parameters. The paper is organized as follows: Section 2 provides background results related to QC codes and their constituent codes. This section also includes the foundations of constructing QECCs from classical ones. In Section 3, we present the main results of this paper, namely, the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (or self-dual) with respect to the Euclidean inner product. We show that the self-orthogonal QC codes in this construction satisfy the square-root-like lower bound to their minimum distance. We also introduce the notion of Galois-closed codes to obtain QC codes that are self-orthogonal with respect to the Hermitian inner product. Lastly in Section 6, we show how some codes from our construction are used to obtain quantum error-correcting codes with the best known parameters so far. All computations throughout this paper have been done using SageMath or Mathematica.
uantum error correction plays a central role in the development of quantum computing architectures. With the rapid advancement of technologies involving quantum computation and communication, there is a growing demand for effective quantum error-correcting codes (QECCs). These codes are essential for protecting quantum information from noise, which can arise from various types of quantum channels. In the celebrated work of Calderbank, Shor, and Steane Calderbank-Shor1996; Steane1996, the authors proposed the CSS construction that utilizes the duality of certain classical error-correcting codes to construct quantum error-correcting codes. This method is commonly used to construct QECCs over finite fields Ketkar-Klappenecker-Kumar-Sarvepalli2006. QECCs from the CSS construction are most often constructed using self-orthogonal error-correcting codes and its dual, where self-orthogonal codes have applications in lattice theory W98, row-self-orthogonal matrices Massey92; Massey98, linear complementary dual codes Massey98, and many others. Beyond self-orthogonality, the notions of dual-containing and self-dual codes are also significant. Self-dual codes notably exhibit deep connections with group theory Hoehn00; JLLX13, lattices SloaneChapter99, and design theory Sole90, and are recognized for being asymptotically good MACWILLIAMS1972153. Because of these applications, it is a worthwhile endeavor to determine the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes.
The focus of this paper is on constructing quasi-cyclic (QC) error-correcting codes that are self-orthogonal and self-dual under the Euclidean and Hermitian inner products. Quasi-cyclic codes are a natural generalization of cyclic codes and are well-studied in classical coding theory literature. Quasi-cyclic codes have a rich algebraic structure seguin; cem; lingsole and have strong connections with convolutional codes SolomonTilborg. Moreover, there exist QC codes that are asymptotically good longcodes. In 2022, the authors in RB found an upper bound for the locality of QC locally recoverable codes using generator matrices in systematic form where the first block is the identity matrix and the remaining blocks are comprised of smaller circulant matrices. More recently in  BastosAlvarezFloresSalerno, a more concise way of finding an upper bound on the locality of these codes was studied, and a construction of optimal locally recoverable QC codes was presented.
Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in Abdukhalikov-Bag-Panario2023; GalindoHernando; benjwal, the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. Consequently, as these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
We then describe how the codes in our construction yield the existence of QECCs with good parameters. The paper is organized as follows: Section 2 provides background results related to QC codes, their constituent codes, and their different representations that will be used throughout this paper. This section also includes the foundations of constructing QECCs from classical ones. In Section 3, we present the main results of this paper, namely, the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (or self-dual) with respect to the Euclidean inner product. We show that the self-orthogonal QC codes in this construction satisfy the square-root-like lower bound to their minimum distance. We also introduce the notion of Galois-closed codes to obtain QC codes that are self-orthogonal with respect to the Hermitian inner product. Lastly in Section 6, we show how some codes from our construction are used to obtain quantum error-correcting codes. All computations throughout this paper have been done using SageMath or Mathematica.
2 Preliminaries
@@
$d\leq n-k+1$ (1)
and the code is called a maximum distance separable code, or an MDS code, if equality in (1) is achieved. A quintessential example of a maximum distance separable code is the Reed-Solomon code described in RS1958. More generally, if we choose nonzero elements $v_{1},...,v_{n}\in\mathbb{F}_{q}$ and distinct elements $\alpha_{1},...,\alpha_{n}\in\mathbb{F}_{q}$, we define the Generalized Reed-Solomon (GRS) Codes as
$GRS_{n,k}(\mathbf{\alpha},\mathbf{v})=\{(v_{1}f(\alpha_{1}),...,v_{n}f(\alpha_{n}))\mid f(x)\in\mathbb{F}_{q}[x]_{k}\}$ (2)
@@
$\textbf{c}=\begin{pmatrix}c_{00}&c_{01}&...&c_{0,\ell-1}\\ c_{10}&c_{11}&...&c_{1,\ell-1}\\ \vdots&\vdots&\ddots&\vdots\\ c_{m-1,0}&c_{m-1,1}&...&c_{m-1,\ell-1}\end{pmatrix}$ (3)
where the action of the operator $T^{\ell}$ is equivalent to a row shift in (3).
To describe the constituent code representation of a QC code in lingsole, we consider the ring $R_{q,m}^{\ell}:=\left(\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle\right)^{\ell}$, where $\operatorname{gcd}(m,q)=1$. To any codeword c as in (3), we associate the following element of $R_{q,m}^{\ell}$:
$\vec{\textbf{c}}(x):=\left(c_{0}(x),c_{1}(x),...,c_{\ell-1}(x)\right)\in R_{q,m}^{\ell}$ (4)
@@
$c_{j}(x):=c_{0j}+c_{1j}x+c_{2j}x^{2}+\cdots+c_{m-1,j}x^{m-1}\in R_{q,m}.$
By Lemma 3.1 in lingsole, the map $\phi:\mathbb{F}_{q}^{m\ell}\to R_{q,m}^{\ell}$ defined by $\phi(\textbf{c})=\vec{\textbf{c}}(x)$ is an $R_{q,m}$-module isomorphism. Hence, any QC code over $\mathbb{F}_{q}$ of length $m\ell$ and index $\ell$ can be viewed as an $R_{q,m}$-submodule of $R_{q,m}^{\ell}$.
grayThroughout this paper, we will move freely between the representations in (LABEL:flattened_codeword) and (4).
Lemma 1
(lingsole, Lemma 3.1) The map induces a one-to-one correspondence between quasi-cyclic codes over $\mathbb{F}_{q}$ of index $\ell$ and length $\ell m$ and linear codes over $R$ of length $\ell$. AA: The statement of this lemma is redundant from the paragraph directly above. Remove this lemma?G.: Yes. In fact we cite this lemma in the paragraph above
Additionally, given a polynomial $h(x)=h_{n}x^{n}+h_{n-1}x^{n-1}+\cdots+h_{0}\in\mathbb{F}_{q}[x]$, its reciprocal polynomial $h^{*}(x)$ is given by $h^{*}(x)=x^{n}h(x^{-1})$. As seen in lingsole, we may decompose $x^{m}-1\in\mathbb{F}_{q}[x]$ into irreducible factors
$x^{m}-1=\delta\prod_{j=1}^{t}g_{j}(x)g_{j}^{*}(x)\prod_{i=1}^{s}f_{i}(x),$ (5)
where $\delta$ is a non-zero element in $\mathbb{F}_{q}$, $f_{i}(x)$ is a self-reciprocal polynomial for each $1\leq i\leq s$, and $g_{i}^{*}(x)$ is the reciprocal polynomial of $g_{i}(x)$, where $g_{i}(x)\neq g_{i}^{*}(x)$, for each $1\leq j\leq t$. With this decomposition and via the Chinese remainder theorem (CRT), the ring $R_{q,m}$ may be written as
$R_{q,m}=\frac{\mathbb{F}_{q}[x]}{\langle x^{m}-1\rangle}\cong\Bigg(\bigoplus_{j=1}^{t}\bigg(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\bigg)\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg),$ (6)
where we have suppressed the explicit dependence on $x$ in the $g_{j}$ and $f_{i}$ to simplify notation. Let $\alpha$ be an $m$-th primitive root of the unity in an extension of $\mathbb{F}_{q}$ and let
$F_{i}:=\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\cong\mathbb{F}_{q}(\alpha^{u_{i}}),\quad G^{\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\cong\mathbb{F}_{q}(\alpha^{v_{j}}),\quad\text{and}\quad G^{\prime\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\cong\mathbb{F}_{q}(\alpha^{v^{*}_{j}}).$
@@
$R_{q,m}^{\ell}$ $\cong\Bigg(\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}\Bigg)\oplus\left(\bigoplus_{i=1}^{s}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{\ell}\right)$ (7)
$\cong\Bigg(\bigoplus_{j=1}^{t}(G^{\prime}_{j}\oplus G^{\prime\prime}_{j})^{\ell}\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}F_{i}^{\ell}\bigg).$ (8)
@@
Define codes $C_{i}\subseteq F_{i}^{\ell}$, $C^{\prime}_{j}\subseteq G_{j}^{{}^{\prime}\ell}$, and $C^{\prime\prime}_{j}\subseteq G_{j}^{{}^{\prime\prime}\ell}$ for $1\leq i\leq s$ and $1\leq j\leq t$. We call these codes the constituent codes of the QC code $\mathcal{C}$ which has the following CRT decomposition:
$\mathcal{C}\cong\Bigg(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus C_{j}^{\prime\prime})\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}C_{i}\bigg).$ (9)
@@
The following lemma gives an easy way to compute the dimension of a QC code $\mathcal{C}$ from its constituent codes. We slightly modified the notation of the following result from BastosAlvarezFloresSalerno to align with the notation in this paper.
Lemma 2
@@
(BastosAlvarezFloresSalerno, Lemma 2) Let $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ be an $m\ell,k]$ QC code and let $C_{\mu}\subseteq\left(\frac{\mathbb{F}_{q}[x]}{\langle b_{j}\rangle}\right)^{\ell}$ be its respective constituent $\ell,k_{\mu}]$ codes. If the nontrivial constituent codes of $\mathcal{C}$ are given by $C_{1},\ldots,C_{h}$, then $k=\sum_{\mu=1}^{h}k_{\mu}\deg(b_{\mu}).$
Moreover, if the nonzero constituent codes of $\mathcal{C}$ are denoted by $C_{1},C_{2},...,C_{h}$, we assume from now on that their minimum distances satisfy the following descending-chain condition
$d(C_{1})\geq d(C_{2})\geq\cdots\geq d(C_{h}),$ (10)
which will be utilized when computing a lower bound for the minimum distance of $\mathcal{C}$.
gray
2.1 Jensen’s Concatenated Description for QC Codes
et $r\in\{1,$$\ell\}$ and let $\mathcal{C}\subseteq R_{q,m}^{\ell}$ be a QC code as an $R$-module whose generators are given as
$\left\{\left(a_{1,0}(x),...,a_{1,\ell-1}(x)\right),...,\left(a_{r,0}(x),...,a_{r,\ell-1}(x)\right)\right\}\subset R_{q,m}^{\ell}.$
Then for each i=1,2,,s,i=1,2,...,s, we have G.: $\xi$ or $\alpha$ ? Good catch! Thanks! I was just copying from Jensen’s paper and did not notice the inconsistent notation
$C_{i}=\mathrm{Span}_{F_{i}}\{\left(a_{b,0}(\alpha^{u_{i}}),...,a_{b,\ell-1}(\alpha^{u_{i}})\right)\mid 1\leq b\leq r\}$
where $F_{i}:=\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\cong\mathbb{F}_{q}(\alpha^{u_{i}})$.
uppose that the nontrivial constituent codes of a QC code $\mathcal{C}$ are given by $C_{1},\ldots,C_{h}$. Let $\theta_{i}$ be the generating primitive idempotent for the minimal cyclic code $\langle\theta_{i}\rangle$ in consideration, where details on $\theta_{i}$ can be found in (OO24, Section 2). We then have the following isomorphism given by the following maps
$\phi:\langle\theta_{i}\rangle\to\mathbb{E}_{i},\;a(x)\mapsto a(\xi^{u_{i}})\;\;\;\text{and}\;\;\psi:\mathbb{E}_{i}\to\langle\theta_{o}\rangle,\delta\mapsto\sum_{k=0}^{m-1}a_{k}x^{k}$
where $a_{k}=\frac{1}{m}\mathrm{Tr}_{\mathbb{E}/\mathbb{F}_{q}}\left(\delta\xi^{-ku_{i}}\right)$. If $C_{i}$ is a constituent code of $\mathcal{C}$ for $1\leq i\leq h$, then we denote its concatenation with $\langle\theta_{i}\rangle$ by $\langle\theta_{i}\rangle\square C_{i}$. Jensen jensen provided the following description of QC codes via concatenation, and we will adopt the formulation presented in OO24.
Theorem 3
(OO24, Theorem 2.1)
  • (i)
    Let $\mathcal{C}$ be a QC code of length $m\ell$ and index $\ell$ over $\mathbb{F}_{q}$. Then for some subset $\mathcal{I}$ of $\{1,...,s\}$, there exist linear codes $C_{i}$ of length $\ell$ over $F_{i}$, which can be explicitly described, such that
    $\mathcal{C}\cong\bigoplus_{i\in\mathcal{I}}\langle\theta_{i}\rangle\square C_{i}.$
  • (ii)
    Conversely, let $C_{i}$ be a linear code in $F_{i}^{\ell}$, for each $i\in\mathcal{I}\subseteq\{1,...,s\}$. Then,
    $\mathcal{C}\cong\bigoplus_{i\in\mathcal{I}}\langle\theta_{i}\rangle\square C_{i}$
    is a $q$-ary QC code of length $m\ell$ and index $\ell$.
As noted in (GO13, Theorem 4.1), the constituent codes in (9) are equal to the outer codes (i.e., the $C_{i}$’s) in Theorem 3.
2.2 The Cyclic Code Associated to a QC Code and a Lower Bound for its Minimum Distance